Dfuentes642:

Simplify negative 4 over 3 divided by 2 over negative 6.

Dfuentes642:

The choices are -4 -2 2 4

Dfuentes642:

Can someone pls help me

Dfuentes642:

thank you

happy2:

When dividing fractions, you can switch the places of the numerator and the denominator in one of the fractions. After that, simply multiply the numerators and denominators of both fractions. Finally, simplify the product of the two fractions.

happy2:

\[(-4/3)\div(2/-6)\] \[(-4/3)\times(-6/2)=24/6\] \[24/6=4\]

happy2:

Hope this helps!